OpenStudy (anonymous):

I will give medals and will fan!!!! Factor. 4x2 + 10x A. 2x(2x + 5) B. 2x2(2x + 5) C. 4x(x + 5) D. x(2x + 5)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Factor. 20mn – 15n A. 4m – 3 B. 5m(4n – 3) C. 5n(4m – 3) D. 5mn(4m + 3n)

OpenStudy (solomonzelman):

Guys, not the answer process is important, but the learning process. Guide the person who asks the question this is an openSTUDY.com .. you know what I mean !

OpenStudy (anonymous):

Sorry, I didn't know.

OpenStudy (the_fizicx99):

Anyone want to explain? Find a coefficient gcf, and variable GCF, \(\ 4x^2 + 10x \Rightarrow\) 4 = 2 * 2 <- 10 = 5 * 2 <- And they both have a variable gcf, x. x^2 = {x} * x x = {x} Now just divide 2x from both terms, 2x( 2x + 5)

OpenStudy (solomonzelman):

It's alright, but now you do! I didn't report knowing you are new to the website. @PleaseSomebodyHelpMe

OpenStudy (anonymous):

Factor. 36w3x4 – 9wx2 A. 3wx(12w2x3 – 3x) B. 6wx(6w2x2 – 1.5x) C. 9wx2(4w2x2 – 1) D. 6w(4w2x4 – x2)

OpenStudy (the_fizicx99):

@sammy34daisy It's one question per post.

OpenStudy (solomonzelman):

@sammy34daisy , yes, one question per post please. you can ask as many questions as you want, just that not all in one post.